# Quantum computing and the stable set problem

### v3

@@ 1.4 Outline and terminology
The rest of this paper is organized as follows. In Section 2, we give an overview of related work that aims to solve the stable set problem with quantum annealing. A QUBO formulation for the stable set problem is derived in Section 3. In Section 4, we discuss some aspects and shortcomings of the given QUBO formulation and present our post-processing procedure. For larger instances, we propose to use the CH-partitioning method and develop its modified version in Section 5. In Section 6, we perform extensive computational study and present results for newly proposed approaches. Finally, we conclude with a short discussion of our results in Section 7.
@@ 2 Related Work
Further numerical experiments were conducted in [3], where the authors compared quantum annealing implementation to several classical algorithms, such as simulating annealing, Gurobi, and some third-party clique-finding heuristics. For their tests, they used D-Wave 2X with roughly $1000$ qubits in Chimera graph architecture. The experiments were performed on random graphs with $45$ vertices and edge probabilities ranging from $0.3$ to $0.9$. D-Wave returned solutions of comparable quality to classical methods, but classical solvers were generally faster for small instances. They also tested subgraphs of D-Wave chimera graphs, where D-Wave returned the best solutions for large instances and showed a substantial computing speed-up. For instances that did not fit D-Wave’s architecture, they proposed decomposing the problem into subproblems and combining solutions, which was effective for sparse graphs.
@@
To solve an optimization problem with D-Wave’s QPU, one should formulate it as an Ising model or a QUBO problem. In [21], respective formulations for many NP-complete and NP-hard problems were introduced. Among others, a QUBO formulation for the stable set was given. The given formulation contains two terms, one for the vertices and one for the edges. While the contribution of a vertex in the solution is rewarded, every edge in the solution is penalized. It is suggested in [21] to set penalization for edge to be strictly greater than a reward for a vertex.
@@ 3.1 Formulation of the stable set problem
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (4) with $-e^{T}x^{*}+\beta(x^{*})^{T}Ax^{*}=-e^{T}x^{*}$, and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an optimum solution for (4) and $X^{*}$ the corresponding subset of $V$. If $X^{*}$ is a stable set, then the optimum value of (4) is $e^{T}x^{*}$, hence $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exists $\{v_{i},v_{j}\}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 00, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x}\leq\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (4):
@@
Let $G=(V,E)$ with $V=\{v_{1},v_{2}\}$, $E=\{\{v_{1},v_{2}\}\}$ and assume $\beta<\frac{1}{2}$. For $x=(1,1)^{T}$, we have that $\text{cost}(x,\beta)=-e^{T}x+\beta x^{T}Ax=-2+2\beta<-1$. Therefore, the minimum of $\text{cost}(x,\beta)$ over all binary vectors is smaller than $-1$, and thus $\alpha(G,\beta)>1=\alpha(G)$.
@@
where $x$ is a binary variable. Since $x_{i}^{2}=x_{i}$, we can write $-e^{T}x$ as $x^{T}(-I)x$. Thus, we set $Q=-I+\beta A$ and obtain the following problem formulation:
@@ 3.2 The scaling problem for quantum annealers
However, a quantum annealer does not operate on the whole interval of the given coefficients, but it scales the coefficients to its operating space. Assume a quantum annealer operates on the coefficient interval $[-1,1]$. Since the coefficients of the function $E_{2}$ exceed this interval, the quantum annealer will correct them by multiplying the function by $\frac{1}{5}$. This yields the new corrected function $E_{2}^{\prime}$:
@@
In the given QUBO formulation for the stable set problem (7), we only have one penalty parameter—the parameter $\beta$. Since we are interested in the relative scaling between the constants, we set $\beta$ in such a manner that the absolute values of the coefficients of all monomials are as similar as possible. Since the monomials $x_{i}^{2}$ have the coefficient $-1$, we note that $2\beta$ should be equal or approximately equal to $|-1|$, as this mitigates the scaling problem the most. This implies that $\beta$ should be close to $\frac{1}{2}$.
@@ 4.1 Analysis of solution
We consider the graph $G[X]$. If $|E(G[X])|\geq|X|$, the statement trivially holds since we can take $X^{\prime}$ to be a single vertex. Now assume that $|E(G[X])|<|X|$. Then, iterating over the set $E(G[X])$, we remove for each edge one of its adjacent vertices. Since there are $|E(G[X])|$ edges, we remove at most that many vertices. As a result, we obtain a stable set of size at least $|X|-|E(G[X])|$.
@@
Let $\beta=1$. Then, the penalty component of the QUBO formulation (7) is $x^{T}Ax$. Expanding $x^{T}Ax$, we notice that each quadratic term appears twice according to Observation 1, resulting in a penalty of $2$ for each edge. Let us assume there exists a stable set of size at least of $25$ included in a subset of vertices $X\subseteq V$ such that $|X|=30$ and $|E(G[X])|=5$. Then, $\text{cost}(X,1)=-20$, so stable sets of sizes $21,22,23,24$ with respective costs of $-21,-22,-23,-24$ will be more favorable than the set $X$. Therefore, even if only approximate, the solver will likely not return the set $X$ as the optimum. On the other hand, as indicated in Proposition 5, we can easily extract a stable set of size $25$ from the set $X$, meaning that the cost function with $\beta=1$ is not accurately representing the desired solution quality.
@@
holds, which implies that $solution-1>cost$. Existence of such graph is ensured by Lemma 7. Since the cost function value of a stable set is equal to the cardinality of the stable set itself, this means that some stable set $X^{\prime}$ of cardinality $|X^{\prime}|=solution-1$ has better cost function value than our set $X$, even though we can easily extract a stable set of size $solution$ out of the set $X$. Therefore, the cost function imprecisely evaluates our solution and can potentially favor some inferior solutions. To avoid this issue, $\epsilon$ should be set to 00, ensuring that $solution-cost$ cannot exceed $1$. Consequently, $\beta=\frac{1}{2}$ is the most suitable parameter value. If it is any larger, the best candidate set might not be detected as we have just shown, and if it is smaller, Equation (4) does not hold anymore due to Corollary 4.
@@ 4.2 Analysis of samples
Let $X\subseteq V$ be a solution returned by a quantum annealer. If $X$ is not a stable set, we know from Proposition 5 that we can extract a stable set of size $|X|-|E(G[X])|$. Now assume that $X\subseteq V$ is any sample returned by a quantum annealer. Then, if $X$ is not a stable set, we could again employ the procedure from Proposition 5 and extract a stable set. Nevertheless, we show that for samples that are not stable sets, we can obtain even better estimations of the size of the stable sets that can be extracted.
Let us return to Example 6. For a subset of vertices $X\subseteq V$ with $|X|=30$ and $|E(G[X])|=5$, Proposition 5 assures us that we can always extract a stable set of size at least $25$ from it. However, it may occur, for instance, that one of the vertices has a degree of $5$, and by removing it, we obtain a stable set of size $29$. This property is generalized with the annihilation number of the graph, a graph parameter that is also an uspper bound on the stability number of a graph, as shown in [30].
@@
The bound (9) may not be effective for general graphs. However, we will not apply it to compute upper bounds for general graphs. Instead, we will use it to compute bounds for graphs induced from the samples returned by a quantum annealer. These induced graphs possess the property of being highly sparse. This is due to the annealer’s objective of finding a set $X$ that minimizes the cost function $-|X|+2\beta|E(G[X])|$. Consequently, in most cases, even when $\beta=\frac{1}{2}$, the returned samples will have negative cost since otherwise, even a single vertex, constituting a stable set, would have a lower cost function than the set $X$. Working with sparse graphs implies low vertex degrees, so the aforementioned upper bound (9), which considers vertex degrees, effectively reduces the number of samples that need to be evaluated.
@@ 4.3 Post-processing procedure
We define solver $S(G,\beta)$ as a function which takes as input a graph $G$ and the value of the parameter $\beta$ and returns the array of solutions $[X]=[X_{1},\ldots,X_{k}]$ such that $\text{cost}(X_{1},\beta)\leq\text{cost}(X_{2},\beta)\leq\ldots\leq\text{cost}(X_{k},\beta)$. Furthermore, we define the computed value $\hat{\alpha}(G,\beta,S)$ as:
@@
Let $S=S(G,\beta)$ denote the solver we use to obtain the initial array of solutions $[X]=[X_{1},\ldots,X_{k}]$, and let $S_{post}=S(G[X_{i}],\beta_{post})$ denote the solver for recomputing the stability number of a sample $X_{i}$. The post-processing steps are as follows: first, we take a look at $X_{1}$, which has the lowest cost, and define the current best solution as $\operatorname{\textit{best}}=|X_{1}|-|E(G[X_{1}])|$. If $X_{1}$ is a stable set, then $|E(G[X_{1}])|=0$, and $|X_{1}|=-\text{cost}(X_{1},\beta)$. Otherwise, from Proposition 5, we know that we can extract a stable set of size at least $|X_{1}|-|E(G[X_{1}])|$ from $X_{1}$.
[…]
@@
             $[C_{1},\ldots,C_{l}]\leftarrow connected\_components(G[X_{i}])$
@@ 5.1 CH-partitioning
[…]
@@ 5.2 Simple CH-partitioning
    in the core partitioning, the set of vertices $V$ is partitioned into $n$ nonempty disjoint core sets, $C_{i}=v_{i}$ for all $i\in\{1,\ldots,n\}$,
@@
    the complementary set $H_{i}$ of vertices for each core set $v_{i}$ is defined as $H_{i}:=N(v_{i})\setminus\{v_{1},\ldots,v_{i-1}\}$.
@@
[…]
@@
Let $K$ be a maximum clique of $G=(V,E)$, and let $|K|=k$. Without loss of generality, let $K=\{v_{p_{1}},v_{p_{2}},\ldots,v_{p_{k}}\}$ such that $1\leq p_{1}<p_{2}<\ldots<p_{k}\leq n$. For the sake of simplicity, we introduce also $L=\{v_{1},v_{2},\ldots,v_{p_{1}-1}\}$, if $p_{1}>1$, and $L=\emptyset$, if $p_{1}=1$. Obviously, $K\cap L=\emptyset$. According to Definition 14, every vertex $v_{j}$ of $K$ is equal to the core set $C_{j}$. Selecting the core set containing the vertex $v_{p_{1}}$, we have that $H_{p_{1}}=N(v_{p_{1}})\setminus L$. For any $w\in K,~{}w\neq v_{p_{1}}$, we have $w\in N(v_{p_{1}})\setminus L=H_{p_{1}}$, therefore, $K\subseteq C_{p_{1}}\cup H_{p_{1}}$.
@@ 5.3 Bounding the partitions
As the first step, we use the previously mentioned advantage of the simple CH-partitioning. Recall that the simple CH-partitioning yields altogether $n$ pairs $(C_{i},H_{i})$. Nevertheless, the size of partitions will become smaller as the number of partitions approaches $n$. Now, assume that we have already computed or just estimated the largest stability number for some partition(s). Then, if the upper bound of some other partition is at most the best-computed value, the respective partition cannot contain a larger stable set.
@@
       $H_{i}\leftarrow N_{\overline{G}}(v_{i})\setminus\cup_{k=1}^{i-1}\{v_{k}\}$

### v4

@@ 1.4 Outline and terminology
The rest of this paper is organized as follows. In Section 2, we give an overview of related work that aims to solve the stable set problem with quantum annealing. A QUBO formulation related to the stable set problem is derived in Section 3. In Section 4, we discuss some aspects and shortcomings of the given QUBO formulation and present our post-processing procedure. For larger instances, we propose to use the CH-partitioning method and develop its modified version in Section 5. In Section 6, we perform extensive computational study and present results for newly proposed approaches. Finally, we conclude with a short discussion of our results in Section 7.
@@ 2 Related Work
Further numerical experiments were conducted in [3], where the authors compared quantum annealing implementation to several classical algorithms, such as simulated annealing, Gurobi, and some third-party clique-finding heuristics. For their tests, they used D-Wave 2X with roughly $1000$ qubits in Chimera graph architecture. The experiments were performed on random graphs with $45$ vertices and edge probabilities ranging from $0.3$ to $0.9$. D-Wave returned solutions of comparable quality to classical methods, but classical solvers were generally faster for small instances. They also tested subgraphs of D-Wave chimera graphs, where D-Wave returned the best solutions for large instances and showed a substantial computing speed-up. For instances that did not fit D-Wave’s architecture, they proposed decomposing the problem into subproblems and combining solutions, which was effective for sparse graphs.
@@
To solve an optimization problem with D-Wave’s QPU, one should formulate it as an Ising model or a QUBO problem. In [21], respective formulations for hard problems were introduced. Among others, a QUBO formulation for the stable set was given. The given formulation contains two terms, one for the vertices and one for the edges. While the contribution of a vertex in the solution is rewarded, every edge in the solution is penalized. It is suggested in [21] to set penalization for edge to be strictly greater than a reward for a vertex.
@@ 3.1 Formulation of the stable set problem
Let $x^{*}\in\{0,1\}^{n}$ be an optimal solution for the problem (3). Since the formulation (3) is equivalent to the formulation (2), we have that $x^{*}_{i}x^{*}_{j}=0$ for all $\{v_{i},v_{j}\}\in E$. Hence, the optimal solution $x^{*}$ is a stable set. Furthermore, $x^{*}$ is a feasible solution for the formulation (4) with $e^{T}x^{*}+\beta(x^{*})^{T}Ax^{*}=-e^{T}x^{*}$, and therefore $\alpha(G)\leq\alpha(G,\beta)$.
Now let $x^{*}\in\{0,1\}^{n}$ be an optimum solution for (4) and $X^{*}$ the corresponding subset of $V$. If $X^{*}$ is a stable set, then the optimum value of (4) is $e^{T}x^{*}$, hence $\alpha(G)\geq\alpha(G,\beta)$. Otherwise, there exists $\{v_{i},v_{j}\}\in E$, such that $x^{*}_{i}=x^{*}_{j}=1$. If we define new $\tilde{x}$, which differs from $x^{*}$ in the $j$-th component only, which is set to 0, this implies $e^{T}\tilde{x}=e^{T}x^{*}-1$ and $\beta\tilde{x}^{T}A\tilde{x} \leq$ $\beta(x^{*})^{T}Ax^{*}-2\beta$. Therefore, this change has the following effect on the cost function of (4):
@@
Let $G=(V,E)$ with $V=\{v_{1},v_{2}\}$, $E=\{\{v_{1},v_{2}\}\}$ and assume $\beta<\frac{1}{2}$. For $x=(1,1)^{T}$, we have that $\text{cost}(x,\beta)=-e^{T}x+\beta x^{T}Ax=-2+2\beta<-1$. Therefore, the minimum of $\text{cost}(x,\beta)$ over all binary vectors is smaller than 1-1, and thus $\alpha(G,\beta)>1=\alpha(G)$.
@@
where $x$ is a binary variable. Since $x_{i}^{2}=x_{i}$, we can write $e^{T}x$ as $x^{T}(-I)x$. Thus, we set $Q=-I+\beta A$ and obtain the following problem formulation:
@@ 3.2 The scaling problem for quantum annealers
However, a quantum annealer does not operate on the whole interval of the given coefficients, but it scales the coefficients to its operating space. Assume a quantum annealer operates on the coefficient interval [1,1][-1,1]. Since the coefficients of the function $E_{2}$ exceed this interval, the quantum annealer will correct them by multiplying the function by $\frac{1}{5}$. This yields the new corrected function $E_{2}^{\prime}$:
@@
In the given QUBO (7), we only have one penalty parameter—the parameter $\beta$. Since we are interested in the relative scaling between the constants, we set $\beta$ in such a manner that the absolute values of the coefficients of all monomials are as similar as possible. Since the monomials $x_{i}^{2}$ have the coefficient 1-1, we note that $2\beta$ should be equal or approximately equal to |1||-1|, as this mitigates the scaling problem the most. This implies that $\beta$ should be close to $\frac{1}{2}$.
@@ 4.1 Analysis of solution
We consider the graph $G[X]$. If $|E(G[X])|\geq|X|$, the statement trivially holds since we can take $X^{\prime}$ to be a single vertex. Now assume that $|E(G[X])|<|X|$. Then, iterating over the set $E(G[X])$, we remove for each edge one of its adjacent vertices. Since there are $|E(G[X])|$ edges, we remove at most that many vertices. As a result, we obtain a stable set of size at least |X||E(G[X])||X|-|E(G[X])|.
@@
Let $\beta=1$. Then, the penalty component of the QUBO formulation (7) is $x^{T}Ax$. Expanding $x^{T}Ax$, we notice that each quadratic term appears twice according to Observation 1, resulting in a penalty of $2$ for each edge. Let us assume there exists a stable set of size at least of $25$ included in a subset of vertices $X\subseteq V$ such that $|X|=30$ and $|E(G[X])|=5$. Then, $\text{cost}(X,1)=-20$, so stable sets of sizes $21,22,23,24$ with respective costs of 21,22,23,24-21,-22,-23,-24 will be more favorable than the set $X$. Therefore, even if only approximate, the solver will likely not return the set $X$ as the optimum. On the other hand, as indicated in Proposition 5, we can easily extract a stable set of size $25$ from the set $X$, meaning that the cost function with $\beta=1$ is not accurately representing the desired solution quality.
@@
holds, which implies that solution1>costsolution-1>cost. Existence of such graph is ensured by Lemma 7. Since the cost function value of a stable set is equal to the cardinality of the stable set itself, this means that some stable set $X^{\prime}$ of cardinality $|X^{\prime}|=solution-1$ has better cost function value than our set $X$, even though we can easily extract a stable set of size solutionsolution out of the set $X$. Therefore, the cost function imprecisely evaluates our solution and can potentially favor some inferior solutions. To avoid this issue, $\epsilon$ should be set to 0, ensuring that solutioncostsolution-cost cannot exceed $1$. Consequently, $\beta=\frac{1}{2}$ is the most suitable parameter value. If it is any larger, the best candidate set might not be detected as we have just shown, and if it is smaller, Equation (4) does not hold anymore due to Corollary 4.
@@ 4.2 Analysis of samples
Let $X\subseteq V$ be a solution returned by a quantum annealer. If $X$ is not a stable set, we know from Proposition 5 that we can extract a stable set of size |X||E(G[X])||X|-|E(G[X])|. Now assume that $X\subseteq V$ is any sample returned by a quantum annealer. Then, if $X$ is not a stable set, we could again employ the procedure from Proposition 5 and extract a stable set. Nevertheless, we show that for samples that are not stable sets, we can obtain even better estimations of the size of the stable sets that can be extracted.
Let us return to Example 6. For a subset of vertices $X\subseteq V$ with $|X|=30$ and $|E(G[X])|=5$, Proposition 5 assures us that we can always extract a stable set of size at least $25$ from it. However, it may occur, for instance, that one of the vertices has a degree of $5$, and by removing it, we obtain a stable set of size $29$. This property is generalized with the annihilation number of the graph, a graph parameter that is also an upper bound on the stability number of a graph, as shown in [30].
@@
The bound (9) may not be effective for general graphs. However, we will not apply it to compute upper bounds for general graphs. Instead, we will use it to compute bounds for graphs induced from the samples returned by a quantum annealer. These induced graphs possess the property of being highly sparse. This is due to the annealer’s objective of finding a set $X$ that minimizes the cost function $|X|+2\beta|E(G[X])|$. Consequently, in most cases, even when $\beta=\frac{1}{2}$, the returned samples will have negative cost since otherwise, even a single vertex, constituting a stable set, would have a lower cost function than the set $X$. Working with sparse graphs implies low vertex degrees, so the aforementioned upper bound (9), which considers vertex degrees, effectively reduces the number of samples that need to be evaluated.
@@ 4.3 Post-processing procedure
We define solver $S(G,\beta)$ as a function which takes as input a graph $G$ and the value of the parameter $\beta$ and returns the array of solutions $X]=[X_{1},\ldots,X_{k}]$ such that $\text{cost}(X_{1},\beta)\leq\text{cost}(X_{2},\beta)\leq\ldots\leq\text{cost}(X_{k},\beta)$. Furthermore, we define the computed value $\hat{\alpha}(G,\beta,S)$ as:
@@
Let $S=S(G,\beta)$ denote the solver we use to obtain the initial array of solutions $X]=[X_{1},\ldots,X_{k}]$, and let $S_{post}=S(G[X_{i}],\beta_{post})$ denote the solver for recomputing the stability number of a sample $X_{i}$. The post-processing steps are as follows: first, we take a look at $X_{1}$, which has the lowest cost, and define the current best solution as $\operatorname{\textit{best}}=|X_{1}|-|E(G[X_{1}])|$. If $X_{1}$ is a stable set, then $|E(G[X_{1}])|=0$, and $|X_{1}|=-\text{cost}(X_{1},\beta)$. Otherwise, from Proposition 5, we know that we can extract a stable set of size at least $|X_{1}|-|E(G[X_{1}])|$ from $X_{1}$.
We go through each sample $X_{i}$ for $i=1,\ldots,k$ and check whether $a(G[X_{i}])>\operatorname{\textit{best}}$. If $G[X_{i}]$ is a stable set, then $a(G[X_{i}])\leq\operatorname{\textit{best}}$ trivially holds since samples are sorted by cost, implying that reevaluating the stability number for this sample cannot yield an improved solution. However, if $G[X_{i}]$ is not a stable set but $a(G[X_{i}])>\operatorname{\textit{best}}$, we recompute the solution on the induced graph $G[X_{i}]$ using $S_{post}$. For this purpose, we use Observation 10 and calculate the stability number of the sample $G[X_{i}]$ by considering its connected components. If the new solution surpasses $\operatorname{\textit{best}}$, we update its value. However, if the condition $a(G[X_{i}])>best$ is not fulfilled, and since $a(G[X_{i}])$ is the upper bound for $\alpha(G[X_{i}])$ according to Theorem 9, we know that the sample $X_{i}$ cannot offer a better solution than the one we currently have, so we can skip the recalculation on $X_{i}$.
@@
    $C_{1},\ldots,C_{l}]\leftarrow connected\_components(G[X_{i}])$
@@ 5.1 CH-partitioning
    the sets $\emptyset\neq C_{i}\subset V$ are called core sets and the set $\{C_{i}\;|\;1\leq i\leq s\}$ is called core partitioning since the core sets must partition $V$, i.e.: $\bigcup_{i}C_{i}=V$ and $C_{i}\cap C_{j}=\emptyset$ for all $i\neq j$,
@@ 5.2 Simple CH-partitioning
    the core partitioning contains only the subsets with one element, i.e., the set of vertices $V$ is partitioned into $n$ nonempty disjoint core sets, $C_{i}=\{v_{i}\}$, for all $i\in\{1,\ldots,n\}$,
@@
    the complementary set $H_{i}$ of vertices for each core set $\{v_{i}\}$ is defined as $H_{i}:=N(v_{i})\setminus\{v_{1},\ldots,v_{i}\}$.
@@
Given a simple CH-partitioning $\mathcal{P}_{S}=\{(C_{i},H_{i})\;|\;1\leq i\leq n\}$ of a graph $G$, the size of the maximum clique of $G$ is equal to $\max_{i}\{k_{i}\}$, where $k_{i}$ is the size of a maximum clique of the subgraph of $G$, induced by $C_{i}\cup H_{i}$ .
@@
Let $K$ be a maximum clique of $G=(V,E)$, and let $|K|=k$. Without loss of generality, let $K=\{v_{p_{1}},v_{p_{2}},\ldots,v_{p_{k}}\}$ such that $1\leq p_{1}<p_{2}<\ldots<p_{k}\leq n$. For the sake of simplicity, we introduce also $L=\{v_{1},v_{2},\ldots,v_{p_{1}}\}$. Obviously, $K\cap L=\{v_{p_{1}}\}$. According to Definition 14, every vertex $v_{j}$ of $K$ is equal to the core set $C_{j}$. Selecting the core set containing the vertex $v_{p_{1}}$, we have that $H_{p_{1}}=N(v_{p_{1}})\setminus L$. For any $w\in K,~w\neq v_{p_{1}}$, we have $w\in N(v_{p_{1}})\setminus L=H_{p_{1}}$, therefore, $K\subseteq C_{p_{1}}\cup H_{p_{1}}$. ∎
@@ 5.3 Bounding the partitions
As the first step, we use the previously mentioned advantage of the simple CH-partitioning. Recall that the simple CH-partitioning yields altogether $n$ pairs $(C_{i},H_{i})$. Nevertheless, the size of partitions $(C_{i},H_{i})$ becomes smaller as $i$ approaches $n$. Now, assume that we have already computed or just estimated the largest stability number for some partition(s). Then, if the upper bound of some other partition is at most the best-computed value, the respective partition cannot contain a larger stable set.
@@
 $H_{i}\leftarrow N_{\overline{G}}(v_{i})\setminus\cup_{k=1}^{i}\{v_{k}\}$